OpenStudy (anonymous):

How do you solve this:Which of the following are solutions to the equation below? x^2 + 3x - 18 = 0

OpenStudy (anonymous):

A. -6 B. 18 C. 6 D. -3 E. 3 F. 9

OpenStudy (anonymous):

check all that apply

OpenStudy (hoblos):

you can just try to plug in each answer and see if it makes the equation true or you can factorize the equation and get the roots do you know how to factorize it ?

jhonyy9 (jhonyy9):

can you factorized it ?

OpenStudy (compassionate):

x=-6,3 E

OpenStudy (jpsmarinho):

You can only use bhaskara, and encounter that 3 and -6 is roots of the equation

OpenStudy (anonymous):

teach me,ill write it down.

OpenStudy (jpsmarinho):

You first find the delta value that is : Delta= b^2 - 4*a*c

OpenStudy (anonymous):

x2 + 6x - 3x - 18 x(x+6) - 3(x+6) (x-3)(x+6)

OpenStudy (anonymous):

3 & -6

OpenStudy (hoblos):

simply.. search for 2 numbers having a product of (-18) and sum of 3 they are 6 and -3 so the equation ends up to be (x-3)(x+6) = 0 the roots are x=3 and x=-6

OpenStudy (jpsmarinho):

After, uses in this formula: X= (-b + square root of delta)/(2*a)

OpenStudy (jpsmarinho):

And X= (-b - square root of delta)/(2*a)

OpenStudy (jpsmarinho):

The variables a, b and c is : ax^2 + bx + c

OpenStudy (anonymous):

i see.so how would you do this one..the same way?

OpenStudy (anonymous):

x^2 - 11x + 24 = 0

OpenStudy (jpsmarinho):

8 and 3 . Yeah in the same way that is write above, using factorization or bhaskara

OpenStudy (anonymous):

ok then,thanks. i got it.